You have started a company and are in

lucklong dash—a

venture capitalist has offered to invest. You own 100% of the company with 4.57 million shares. The VC offers $1.18 million for 800,000 new shares.

a. What is the implied price per​ share?

b. What is the​ post-money valuation?

c. What fraction of the firm will you own after the​ investment?

a. What is the implied price per​ share?

The implied price per share will be

​$nothing

per share.  ​(Round to the nearest​ cent.)

b. What is the​ post-money valuation?

The​ post-money valuation will be

​$nothing

million.

​(Round to two decimal​ places.)

c. What fraction of the firm will you own after the​ investment?

Your fractional ownership will be

nothing​%.

​(Round to one decimal​ place.)
Answer #1

a). Implied price per share = Investment / New Shares = $1,180,000 / 800,000 = $1.475, or $1.48

b). Post-money valuation = Implied price per share * Total number of shares

= $1.475 * [4,570,000 + 800,000]

= $1.475 * 5,370,000 = $7,920,750

c). Your fractional ownership = Your shares / New Total Shares

= 4,570,000 / 5,370,000 = 0.8510, or 85.1%